# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 6.2 Computations on random permutations
In this section we check that for $M\in\mathrm{S}_{N}$ uniformly distributed and $Q$ verifying (H:1), (H:2) and (H:3), the pair $(M,Q)$ is $\ell$-tangle free with high probability, provided that the integer $\ell$ is not too large before the dimension.
@@
There exists $C>0$ such that, for any integer $\ell\geq 1$ the pair $(M,Q)$ is $\ell$-tangle free with probability at least $1-C\frac{\ell^{4}(4d)^{3\ell}}{N}$.
@@
We consider $1\leq\ell\leq\frac{N}{2}$. We say that a path $\gamma:=(x_{1},y_{1},\ell_{1},x_{2}...,x_{T},\ell_{T-1},y_{T},\ell_{T},x_{T+1})$ occurs if for all $t\leq T$, we have $M_{x_{t},y_{t}}=1$. If $(M,Q)$ is $\ell$-tangled then there exists $1\leq T\leq\ell$, $\ell_{1},...,\ell_{T}$ integers such that $\sum_{i=1}^{T}\ell_{i}+T-1\leq\ell$ and a path $\gamma:=(x_{1},y_{1},\ell_{1},x_{2},y_{2},...,x_{T},\ell_{T},y_{T},\ell_{T},x_{T+1})$ that occurs such that there exists $1\leq k,i<k^{\prime}\leq T$ such that either
@@
where $(N)_{t}:=N(N-1)(N-2)\cdots(N-t+1)$. Let $\{x_{t}\}_{t}$ be pairwise distinct elements of $[N]$, except possibly $x_{1}=x_{k+1}$ and $x_{i}=x_{T+1}$ for some $1\leq k,i<T\leq\ell$. We count the number of possible $\gamma$ verifying $I_{1}(k,i)$. For $\{x_{t}\}_{t\notin\{k,T+1\}}$ we have $(N)_{T-1}$ possibilities in regard of the condition of being pairwise distinct. For each $(\ell^{\prime}_{t})_{t\leq T^{\prime}}$ such that $\sum_{t}\ell^{\prime}_{t}+T^{\prime}-1\leq\ell$, we have at most $2d^{\tiny{\sum_{t}\ell^{\prime}_{t}}}\leq 2d^{\ell}$ possibilities for $x_{k+1}$ to have $(B_{1}^{\ell_{1}^{\prime}}\cdots B_{T^{\prime}}^{\ell_{T^{\prime}}^{\prime}})_{x_{1}x_{k+1}}>0$ with $B_{t}\in\{Q,Q^{\ast}\}$ and $B_{t+1}=B_{t}^{\ast}$. Besides the number of $T^{\prime}\leq\ell$ and $(\ell_{t}^{\prime})_{t\leq T^{\prime}}$ such that $\sum_{t}\ell^{\prime}_{t}+T^{\prime}-1\leq\ell$ is given by:
$|\bigsqcup_{p=1}^{\ell}\bigsqcup_{T^{\prime}=1}^{p}\{(\ell^{\prime}_{t})_{t\leq T^{\prime}}:\leavevmode\nobreak\ \sum_{t=1}^{T^{\prime}}\ell_{t}^{\prime}+T^{\prime}-1=p\}|=\sum_{p=1}^{\ell}\sum_{T^{\prime}=1}^{p}\binom{p-1}{T^{\prime}-1}\leq 2^{\ell}.$
@@
There exists a constant $c>0$ such that for any $\mathbf{x}:=(x_{1},...,x_{k}),\mathbf{y}:=(y_{1},...,y_{k})\in[N]^{k}$ with $2k\leq\sqrt{N}$ and any $k_{0}\leq k$, we have:
$\left|\mathbb{E}\large(\prod_{t=1}^{k_{0}}\underline{M}_{x_{t}y_{t}}\prod_{t=k_{0}+1}^{k}M_{x_{t}y_{t}})\right|\leq c2^{b}\left(\frac{1}{N}\right)^{a}\left(\frac{3k}{\sqrt{N}}\right)^{a_{1}},$
where $a:=|\mathrm{A}_{\mathbf{x,y}}|$, b is the number of inconsistent arcs of $(\mathbf{x,y})$ and $a_{1}$ is the number of $1\leq t\leq k_{0}$ such that $(x_{t},y_{t})$ is inconsistent and has multiplicity one.
@@ 7 High trace method
7.1 Operator norm of $\underline{A}_{\#}^{(\ell)}$
@@
Let $(Q,M),\leavevmode\nobreak\ \rho_{n}$ be as defined in 2.2 and for #$\#\in\{\emptyset,r\}$ we consider $A_{\#}$ defined by (3) and (4). For any $c>0$ and $1\leq\ell\leq\frac{\ln(N)}{7K^{\prime}_{\#}\ln(d)}$, there exists $C=C(c,\alpha,n,K)>0$ such that with probability at least $1-N^{-c}$,
$\|\underline{A}_{\#}^{(\ell)}\|\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}.$
We give the details of computations in the case where #=#\#$\emptyset$# = ∅. For any integer $m\geq 1$ and considering as convention $x_{2m+1}:=x_{1}$ we have
$\|\underline{A}^{(\ell)}\|^{2m}=\|\underline{A}^{(\ell)}{\underline{A}^{(\ell)}}^{\top}\|^{m}\leq\operatorname{Tr}\{\left(\underline{A}^{(\ell)}{\underline{A}^{(\ell)}}^{\top}\right)^{m}\}$
$=\sum_{x_{1},...,x_{2m}}\prod_{i=1}^{m}(\underline{A}^{(\ell)})_{x_{2i-1},x_{2i}}(\underline{A}^{(\ell)})_{x_{2i+1},x_{2i}}$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\gamma_{2i}\in\mathrm{F}^{\ell}_{x_{2i+1},x_{2i}}}\prod_{t=1}^{T_{2i}}\underline{M}_{x_{2i,t},y_{2i,t}}(Q^{\ell_{2i,t}})_{y_{2i,t}x_{2i,t+1}}\bigg{\}}$
@@
$\|\underline{A}^{(\ell)}\|^{2m}\leq\sum_{\vec{T}\in[\ell]^{2m}}\sum_{\gamma\in\mathrm{W}_{\ell,\vec{T}}}\prod_{i=1}^{m}\prod_{t=1}^{T_{i}}\underline{M}_{x_{i,t}y_{i,t}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (27)
[…]
$x_{i,t+1}=x,\leavevmode\nobreak\ \leavevmode\nobreak\ x_{i^{\prime},t^{\prime}+1}=x^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ y_{i,t}=y_{i^{\prime},t^{\prime}}.$
@@
We also set the set of arcs of $\gamma$, denoted $\mathrm{A}_{\gamma}$ as the set of distinct pairs $(x_{i,t},y_{i,t})$. We then consider the partition $\{\mathrm{W}_{\ell,\vec{T}}(s,a,p)\}_{s,a,p}$ where $\mathrm{W}_{\ell,\vec{T}}(s,a,p)$ is the set of $\gamma\in\mathrm{W}_{\ell,\vec{T}}$ such that $|X_{\gamma}|=s$, $|\mathrm{A}_{\gamma}|=a$ and $|X_{\gamma}/\sim|=|\{\operatorname{cc}(x)\leavevmode\nobreak\ ;\leavevmode\nobreak\ x\in X_{\gamma}\}|=s-p$. We denote $G_{\gamma}=(X_{\gamma}/\sim,\mathrm{A}_{\gamma})$ the graph where for all arc $(x,y)=(x_{i,t},y_{i,t})\in\mathrm{A}_{\gamma}$ we put an edge $(\operatorname{cc}(x_{i,t}),\operatorname{cc}(x_{i,t+1}))$ of color $(x,y)$.
@@
We represented the corresponding graphs $G_{\gamma}$ and $K_{\gamma}$ in Figure 6. We have in this case $X_{\gamma}:=\{2,4,8,7\}$, a unique edge in $K_{\gamma}$ given by $\{4,8\}$ (dashed in Figure 6), $(X_{\gamma}/\sim)=\{\operatorname{cc}_{1},\operatorname{cc}_{2},\operatorname{cc}_{3}\}$ and $\mathrm{A}_{\gamma}=\{(2,1),(2,3),(4,6),(7,3)\}$. One can also refer to [BDH22, Figure 4,5] and [Bor15, Figure 3,4] for representation of other examples of this graphs.
@@
Proceeding similarly in the case where #$\#=r$ and taking the expectation of (27) we have
$\mathbb{E}\|\underline{A}_{\#}^{(\ell)}\|^{2m}\leq\sum_{s,a,p,\vec{T}}\sum_{\leavevmode\nobreak\ \gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}\mu(\gamma)q_{\#}(\gamma),$
@@
$\leavevmode\nobreak\ q(\gamma):=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ q_{r}(\gamma):=(1-r)^{T}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(\{rQ\}^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (29)
@@
$\mathbb{E}\|\underline{A}_{\#}^{(\ell)}\|^{2m}\leq\sum_{s,a,p,\vec{T}}|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}{\max}\left\{|\mu(\gamma)|\sum_{\begin{subarray}{c}\gamma^{\prime}\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)\\ \gamma^{\prime}\sim\gamma\end{subarray}}q_{\#}(\gamma^{\prime})\right\}.$ (30)
@@
The following section is dedicated to the proof of 7.2.

### v4

@@ 7.3 Computations on random permutations
In this section we check that for $M\in\mathrm{S}_{N}$ uniformly distributed and $Q$ deterministic verifying (H:3), the pair $(M,Q)$ is $\ell$-tangle free with high probability, provided that the integer $\ell$ is not too large before the dimension. Recall that for $Q$ a matrix we denote
$d:=\underset{x\in[N]}{\max}|\{y\in[N]\penalty 10000\ ;\penalty 10000\ Q_{xy}\neq 0\}|\vee\underset{x\in[N]}{\max}|\{y\in[N]\penalty 10000\ ;\penalty 10000\ Q^{\ast}_{xy}\neq 0\}|.$
@@
Let $Q$ be a deterministic matrix and $M\sim\mathcal{U}(\mathrm{S}_{N})$. There exists $C>0$ such that, for any integer $\ell\geq h$, the pair $(M,Q)$ is $\ell$-tangle free with probability at least $1-C\frac{\ell^{4}(4d)^{3\ell}}{N}$.
@@
We consider $1\leq\ell\leq\frac{N}{2}$. We say that a path $\gamma:=(x_{1},y_{1},\ell_{1},x_{2}...,x_{T},\ell_{T-1},y_{T},\ell_{T},x_{T+1})$ occurs if for all $t\leq T$, we have $M_{x_{t},y_{t}}=1$. If $(M,Q)$ is $\ell$-tangled then there exists $1\leq T\leq\ell$, $\ell_{1},...,\ell_{T}$ integers verifying $\sum_{i=1}^{T}\ell_{i}+T-1=\ell$ and a path $\gamma:=(x_{1},y_{1},\ell_{1},x_{2},y_{2},...,x_{T},\ell_{T},y_{T},\ell_{T},x_{T+1})$ such that there exists $1\leq k,i<k^{\prime}\leq T$ such that either
@@
where $(N)_{t}:=N(N-1)(N-2)\cdots(N-t+1)$. Let $\{x_{t}\}_{t}$ be pairwise distinct elements of $[N]$, except possibly $x_{1}=x_{k+1}$ and $x_{i}=x_{T+1}$ for some $1\leq k,i<T\leq\ell$. We count the number of possible $\gamma$ verifying $I_{1}(k,i)$. For $\{x_{t}\}_{t\notin\{k,T+1\}}$ we have $(N)_{T-1}$ possibilities in regard of the condition of being pairwise distinct. For each $(\ell^{\prime}_{t})_{t\leq T^{\prime}}$ such that $\sum_{t}\ell^{\prime}_{t}+T^{\prime}-1\leq\ell$, we have at most $2d^{\tiny{\sum_{t}\ell^{\prime}_{t}}}\leq 2d^{\ell}$ possibilities for $x_{k+1}$ to have $(B_{1}^{\ell_{1}^{\prime}}\cdots B_{T^{\prime}}^{\ell_{T^{\prime}}^{\prime}})_{x_{1}x_{k+1}}>0$ with $B_{t}\in\{Q,Q^{\ast}\}$ and $B_{t+1}=B_{t}^{\ast}$. Besides the number of $T^{\prime}\leq\ell$ and $(\ell_{t}^{\prime})_{t\leq T^{\prime}}$ such that $\sum_{t}\ell^{\prime}_{t}+T^{\prime}-1\leq\ell$ is given by
$|\bigsqcup_{p=1}^{\ell}\bigsqcup_{T^{\prime}=1}^{p}\{(\ell^{\prime}_{t})_{t\leq T^{\prime}}:\penalty 10000\ \sum_{t=1}^{T^{\prime}}\ell_{t}^{\prime}+T^{\prime}-1=p\}|=\sum_{p=1}^{\ell}\sum_{T^{\prime}=1}^{p}\binom{p-1}{T^{\prime}-1}\leq 2^{\ell}.$
@@
There exists a constant $c>0$ such that for any $\mathbf{x}:=(x_{1},...,x_{k}),\mathbf{y}:=(y_{1},...,y_{k})\in[N]^{k}$ with $2k\leq\sqrt{N}$ and any $k_{0}\leq k$, we have
$\left|\mathbb{E}\left(\prod_{t=1}^{k_{0}}\underline{M}_{x_{t}y_{t}}\prod_{t=k_{0}+1}^{k}M_{x_{t}y_{t}}\right)\right|\leq c2^{b}\left(\frac{1}{N}\right)^{a}\left(\frac{3k}{\sqrt{N}}\right)^{a_{1}},$
where $a:=|\mathrm{A}_{\mathbf{x,y}}|$, b is the number of inconsistent arcs of $(\mathbf{x,y})$ and $a_{1}$ is the number of $1\leq t\leq k_{0}$ such that $(x_{t},y_{t})$ is consistent and has multiplicity one.
@@ 8 High trace method
8.1 Operator norm of $\underline{A}_{\mathrm{x}}^{(\ell)}$
@@
Let $(Q,M),\penalty 10000\ \rho_{n,\mathrm{x}}$ be as defined in ˜2.1 and for $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$ we consider $A_{\mathrm{x}}$ defined by (4) and (5). For any $c>0$ there exists $C=C(c,\alpha,n,K)>0$ such that with probability at least $1-N^{-c}$,
$\|\underline{A}_{\mathrm{x}}^{(\ell)}\|\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}$
for any $1\leq\ell\leq\frac{\ln(N)}{7K_{\mathrm{x}}(\ln(d)\vee 2\ln(\delta\vee 1))}$ where $K_{\mathrm{x}}$ is defined by (20).
We give the details of computations in the case where $\mathrm{x}=\mathrm{s}$. For any integer $m\geq 1$ and considering as convention $x_{2m+1}:=x_{1}$ we have
$\|\underline{A}_{\mathrm{s}}^{(\ell)}\|^{2m}=\|\underline{A}_{\mathrm{s}}^{(\ell)}{\underline{A}_{\mathrm{s}}^{(\ell)}}^{\top}\|^{m}\leq\operatorname{Tr}\{\left(\underline{A}_{\mathrm{s}}^{(\ell)}{\underline{A}_{\mathrm{s}}^{(\ell)}}^{\top}\right)^{m}\}$
$=\sum_{x_{1},...,x_{2m}}\prod_{i=1}^{m}(\underline{A}_{\mathrm{s}}^{(\ell)})_{x_{2i-1},x_{2i}}(\underline{A}_{\mathrm{s}}^{(\ell)})_{x_{2i+1},x_{2i}}$
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{\gamma_{2i}\in\mathrm{F}^{\ell}_{x_{2i+1},x_{2i}}}\prod_{t=1}^{T_{2i}}\underline{M}_{x_{2i,t},y_{2i,t}}(Q^{\ell_{2i,t}})_{y_{2i,t}x_{2i,t+1}}\bigg\}$
@@
$\|\underline{A}_{\mathrm{s}}^{(\ell)}\|^{2m}\leq\sum_{\vec{T}\in[\ell]^{2m}}\sum_{\gamma\in\mathrm{W}_{\ell,\vec{T}}}\prod_{i=1}^{m}\prod_{t=1}^{T_{i}}\underline{M}_{x_{i,t}y_{i,t}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (30)
We want to estimate the expectation of the above sum using Section˜7.3 and the fact that $\gamma$ is composed by $2m$ $\ell$-tangled free paths. We decompose $\mathrm{W}_{\ell,\vec{T}}$ using the size of the support for $\gamma$. For $\gamma=(\gamma_{1},...,\gamma_{2m})\in\mathrm{W}_{\ell,\vec{T}}$, we set $X_{\gamma}:=\{x_{i,t}\penalty 10000\ ;i\in[2m],\penalty 10000\ t\in[T_{i}+1]\}$ and $Y_{\gamma}:=\{y_{i,t}\penalty 10000\ ;i\in[2m]\penalty 10000\ ,t\in[T_{i}]\}$. We consider the graph $K_{\gamma}$ defined as follows. The vertex set is given by $X_{\gamma}$ and for any $^{\prime}\in X_{\gamma}, {x,x}\{x,x^{\prime}\}$ is a vertex if there exists times $(i,t)$ and $(i^{\prime},t^{\prime})$ such that $i,i^{\prime}\in[2m]$ and $(t,t^{\prime})\in[T_{i}]\times[T_{i^{\prime}}]$ such that
$x_{i,t+1}=x,\penalty 10000\ \penalty 10000\ x_{i^{\prime},t^{\prime}+1}=x^{\prime}\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ y_{i,t}=y_{i^{\prime},t^{\prime}}.$
@@
We also set the set of arcs of $\gamma$, denoted $\mathrm{A}_{\gamma}$ as the set of distinct pairs $(x_{i,t},y_{i,t})$. We then consider the partition $\{\mathrm{W}_{\ell,\vec{T}}(s,a,p)\}_{s,a,p}$ where $\mathrm{W}_{\ell,\vec{T}}(s,a,p)$ is the set of $\gamma\in\mathrm{W}_{\ell,\vec{T}}$ such that $|X_{\gamma}|=s$, $|\mathrm{A}_{\gamma}|=a$ and $|X_{\gamma}/\sim|=|\{\operatorname{cc}(x)\penalty 10000\ ;\penalty 10000\ x\in X_{\gamma}\}|=s-p$. We denote $G_{\gamma}=(X_{\gamma}/\sim,\mathrm{A}_{\gamma})$ the graph where for all arc $(x,y)=(x_{i,t},y_{i,t})\in\mathrm{A}_{\gamma}$ we put an edge $(\operatorname{cc}(x_{i,t}),\operatorname{cc}(x_{i,t+1}))$ of color $(x,y)$.
@@
We represented the corresponding graphs $G_{\gamma}$ and $K_{\gamma}$ in Figure˜6. We have in this case $X_{\gamma}:=\{2,4,8,7\}$, a unique edge in $K_{\gamma}$ given by $\{4,8\}$ (dashed in Figure˜6), $(X_{\gamma}/\sim)=\{\operatorname{cc}_{1},\operatorname{cc}_{2},\operatorname{cc}_{3}\}$ and $\mathrm{A}_{\gamma}=\{(2,1),(2,3),(4,6),(7,3)\}$. One can also refer to [BDH22, Figure 4,5] and [BOR15, Figure 3,4] for representation of other examples of this graphs.
$\operatorname{cc}_{1}$$\operatorname{cc}_{2}$$\operatorname{cc}_{3}$2143876(2,3)(2,3)(4,6)(4,6)(2,1)(2,1)$(7,3)$
@@
Proceeding similarly in the case where $\mathrm{x}=\mathrm{c}$ and taking the expectation of (30) we have
$\mathbb{E}\|\underline{A}_{\mathrm{x}}^{(\ell)}\|^{2m}\leq\sum_{s,a,p,\vec{T}}\sum_{\penalty 10000\ \gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}\mu(\gamma)q_{\mathrm{x}}(\gamma),$
@@
$\penalty 10000\ q_{\mathrm{s}}(\gamma):=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ q_{\mathrm{c}}(\gamma):=(1-r)^{T}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(\{rQ\}^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (32)
@@
$\mathbb{E}\|\underline{A}_{\mathrm{x}}^{(\ell)}\|^{2m}\leq\sum_{s,a,p,\vec{T}}|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}{\max}\left\{|\mu(\gamma)|\sum_{\begin{subarray}{c}\gamma^{\prime}\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)\\ \gamma^{\prime}\sim\gamma\end{subarray}}q_{\mathrm{x}}(\gamma^{\prime})\right\}.$ (33)
@@
The following section is dedicated to the proof of ˜8.1.